OpenStudy (anonymous):

Using dimensional analysis, make the following conversions. Show all steps a)1.0 m^3 to ml b)1.0 nm to km Give an example of an entity that is an atom, a formula unit and a molecule. Use examples Using nitrogen gas as an example, show the relationship between moles, mass and number of entities

OpenStudy (anonymous):

1. Welcome to OpenStudy! :-) Most under-read thing on the site for new people (including myself initially: http://openstudy.com/code-of-conduct) 2. One question per topic is preferred 3. Dimensional analysis, if I recall is just a fancy way of saying convert units like so:  Important note: (100 cm)\(^3\) = 100\(^3\) cm\(^3\) = 1000000 cm\(^3\) \[1 \cancel{m^3} \cdot \left[ \frac{(100\ \cancel{cm)^3}}{(1\ \cancel{m)^3}} \cdot \frac{1\ mL}{1\ \cancel{cm^3}} \right]=?\] \[1\ \cancel{nm} \cdot \left[ \frac{1\ \cancel{m}}{1,000,000,000\ \cancel{nm}} \cdot \frac{1\ km}{1000\ \cancel{m}} \right]=?\] 4. Always write your units, so you can see if things cancel out as they should to get the right answer. This remains true in upper level chemistry and physics ;-)